OpenStudy (anonymous):

Express y as a function of x. The constant C is a positive number. ln(y-1)=-6x+lnC Y=__________

terenzreignz (terenzreignz):

I'm pretty sure there's an =sign somewhere...

OpenStudy (anonymous):

sorry just fixed it! thank you for catching it!

terenzreignz (terenzreignz):

Well you want to make it so that y is alone, and, you know... *alone* on one side of the equation. First you'll want to get rid of that nasty "ln" that still tails it... any ideas? :)

terenzreignz (terenzreignz):

Here's a tip.. \[\huge a=b \ \ \rightarrow \ \ e^a=e^b\]

OpenStudy (anonymous):

umm instead of ln you would use e ?

terenzreignz (terenzreignz):

No... but we would raise e to both sides of the equation :) \[\huge \ln(y-1)=-6x+\ln C\] \[\huge e^{\ln(y-1)}=e^{-6x+\ln C}\]

OpenStudy (anonymous):

Ohhh! then how do we get rid of ln from that?

terenzreignz (terenzreignz):

Well, you had best make use of this property \[\huge e^{\ln b}=b\]

terenzreignz (terenzreignz):

That is to say, raising e to a power, and ln cancel each other out.

OpenStudy (anonymous):

okay?

terenzreignz (terenzreignz):

Well? What does that mean for \[\huge e^{\ln(y-1)}=?\]

OpenStudy (anonymous):

will equal to ln(y-1)

OpenStudy (anonymous):

I mean y-1

terenzreignz (terenzreignz):

That's right :) So it becomes \[\huge y -1 = e^{-6x+\ln C}\] Now, I want you to pay attention to the right side... What do the laws of exponents say about \[\huge a^{m+n=?}\]

terenzreignz (terenzreignz):

\[\huge a^{m+n}=a^ma^n\] right?

OpenStudy (anonymous):

yes!

terenzreignz (terenzreignz):

So, the right side may be written \[\huge y-1=e^{-6x}e^{\ln C}\]

terenzreignz (terenzreignz):

And now we have \(\Large e^{\ln C}\) What is this equal to?

OpenStudy (anonymous):

sorry I have no clue now.

terenzreignz (terenzreignz):

Same with what you did for \(\Large e^{\ln(y-1)}\)

OpenStudy (anonymous):

it will equal to C

terenzreignz (terenzreignz):

That's right :) So finally \[\huge y-1 =Ce^{-6x}\] And now, you just have to bring the 1 to the right side, and then y will be alone, as a function of x ^.^

OpenStudy (anonymous):

so it would be Y= Ce^-6x +1

terenzreignz (terenzreignz):

Yes it will :)

terenzreignz (terenzreignz):

sorry, I was looking at another question :)

OpenStudy (anonymous):

haha its okay thank you a lot!